# Geodesic structure of spacetime near singularities

### v1

@@ I Introduction
An important first step in understanding the role of quantum effects in regions of strong gravitational fields is to understand the behavior of geodesic flows - behavior of geodesics emanating from a single point - in vicinity of the singularity. This is so because many important features of classical and quantum dynamics depend on this behavior; see, for instance, Gutzwiller (1990) for an extensive textbook discussion on this point. While the structure of these flows is well understood in geodesically convex neighborhoods of regular spacetime points, not much is known about their behavior in regions containing curvature singularities. This is to be expected since covariant Taylor expansions that work well in the former case breaks down in the latter, and hence the analysis of geodesic flows directly in terms of curvature related quantities can not be done for generic singularities. We derive here, for the first time, explicit behavior of Synge’s world function $\Omega(x,y)$ and van Vleck determinant $\Delta(x,y)$ - two important bi-scalars characterizing geodesic flows - in the vicinity of FLRW and Kasner like singularities. In the process, we improve upon an old result by Buchdahl Buchdahl and Warner (1980) who calculated the series expansion of $\sqrt{|\Omega|}$” in FLRW and obtained a series whose terms can be shown to diverge as one approaches the singularity at $t\to 0$. This obviously un-physical result, as we shall see, is an artifact of expanding a wrong function (and manifests, in fact, even in the Minkowski limit!), and was presumably not noticed by Buchdahl. Here, we focus instead on $\Omega(x,y)$ itself, which we take to be the only meaningful quantity.
@@ I.1 Synge’s world function and van Vleck determinant
Synge’s world function and van Vleck determinant are two important bi-scalars that characterize measurements in curved spacetimes, and also appear ubiquitously in quantum field theory. Two-point (and higher) correlators of quantum fields in curved spacetime have a universal expansion in the world function near the coincidence limit. These expansions form the backbone of the point-splitting program for regularization in curved spacetime, and has been used developed and studied extensively by de Witt, Cristensen, Parker and others in deriving covariant expansions of effective lagrangians in curved spacetime, covariant formulation of path integral etc. The use of the world function as a fundamental object to study spacetime dates back to Synge, where it is defined operationally as half the square of the geodesic distance between the two spacetime points (denoted by $\Omega(x,x^{\prime})$). The spacetime metric $g_{ab}(x)$ can be constructed as the coincidence limit of $\nabla_{a}\nabla_{b}\Omega(x,y)$, i.e.
@@
Similarly, higher order coincidences gives us information about the spacetime curvature. Thus world function can be used to study tensor expansions on a given spacetime manifold in a covariant manner. The only hindrance to it’s use (apart from existence of multiple geodesics connecting two points) is the non-existence of any closed-form expression for geodesic interval as a function of spacetime coordinates even in highly symmetric spacetimes. In fact, to the best of our knowledge, closed form expressions are known only in maximally symmetric and plane-wave spacetimes Hari and Kothawala (2021),Harte and Drivas (2012). It would surprise the in-cognoscenti that the simplest and most widely encountered spacetimes such as Schwarzschild and FLRW have no closed form expression for $\Omega$. In this work, we derive a systematic representation for the world function in FLRW and Bianchi spacetimes, in a form that is applicable near singularities in these spacetimes.
@@
van Vleck determinant is a bivector of central importance that can be computed using the world function directly. It appears at various places, like fluctuation factor in semi-classical approximation to path integrals in curved spacetime, reconstruction of spacetime while defining q-metric using van Vleck determinant, heat kernel expansions in curved spacetime, etc. Given a geodesic with endpoints $x$ and $z$, It is defined as,
@@
Where $\Omega(x,z)$ is Synge’s world function. More intuitively, it’s a quantitative measure of density of geodesics emanating from a given point as shown in fig.1.
@@ I.2 Summary of results
The table 1 gives the summary of the coincidence and singularity limits obtained for the various biscalars in FLRW and Bianchi Type I spacetime. We obtained the closed-form expression for Synge’s world function for FLRW and Kasner (Schwarzschild) spacetime in the form of an infinite series. The limits of $\Delta$ and $\Box\Delta^{1/2}$ and $K=\textrm{Tr}(K_{ab})$ are computed to study the geodesic and causal structure near the curvature singularity. The lightcones structure is illustrated in fig. 3 and fig. 4 by plotting null geodesics along with $\Omega=$ constant surfaces about different points in the neighbourhood of FLRW and Kasner (Schwarzschild) singularity.
@@
$\Delta(t,T)$ $\textrm{matter}:1+\frac{\varepsilon^{2}}{9T^{2}}+\cdots$ $\textrm{radiation}:1+\frac{\varepsilon^{2}}{8T^{2}}+\cdots$ $\textrm{matter}:\frac{t}{27T}$ $\textrm{radiation}:\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}$ $1+O(\varepsilon^{4})$ $\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}+\cdots$
$\Box\Delta^{1/2}(t,T)$ $K(t,T)$ $\textrm{matter}:\frac{2}{9T^{2}}$ $\textrm{radiation}:\ \ 0$ $\frac{3}{\varepsilon}+\frac{\ddot{a}}{a}\varepsilon$ $\textrm{matter}:\frac{2267}{2187\sqrt{21}t^{2}}$ $\textrm{radiation}:\ \ \ 0$ $\frac{3q}{t}$ $\frac{3}{\varepsilon}-\frac{8\varepsilon^{3}}{1215T^{4}}+$... $\frac{8}{3t}$
@@ II.1 FLRW spacetime
The (9) is a non-perturbative closed-form expression in terms of the geodesic length $\ell$ in geodesic distance on the three-space of constant curvature. The complication of further simplification arises because of the time-dependent scale factor $a(t)$, which is kept open for generality. As another self-consistency check, it can be verified that (IV) satisfies the HJ equation up to order $\mathcal{O}(\ell^{4})$ with;
@@
$\Delta(t,T)\simeq\frac{1}{a(t)^{3}a(T)^{3}F(t,T)^{3}}+\frac{20F(t,T)^{3}Q(t,T)-F^{(1,1)}(t,T)}{2a(t)^{3}a(T)^{3}F(t,T)^{5}}\ell^{2}+\mathcal{O}(\ell^{4})$ (11)
@@
where $I_{n}(a_{i})$ is defined by (31). We now consider the limiting behaviour of $\Delta(t,T)$ along a timelike geodesic in coincidence and singularity limit in matter and radiation dominated FLRW spacetime. The expansion in $\epsilon=T-t<<T$ (coincidence limit) and (Singularity limit) $T<<t,T\to 0$ are obtained as follows,
@@
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $\simeq$ $1+\frac{\epsilon^{2}}{9T^{2}}+\mathcal{O}(\epsilon^{3})$ (15)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $\simeq$ $\frac{t}{27T}+$... (16)
@@
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $\simeq$ $1+\frac{\epsilon^{2}}{8T^{2}}+\mathcal{O}(\epsilon^{3})$ (17)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $\simeq$ $\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}+$... (18)
@@ II.2 Bianchi Type I singularities
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are integration constants. The eqn. (25) can be inverted using Lagrange-Good formula Abdesselam (2003) for multivariables to give a series expansion in (xX),(yY),(zZ)(x-X),(y-Y),(z-Z), where
@@
where $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $a,b=1,2,3$. The $\mathcal{O}((x_{i}-X_{i})^{4})$ expansion for $\Omega$ using (27) gives,
@@
Where $\ell^{2}_{\perp}=(x-X)^{2}+(y-Y)^{2}$. It is easy to verify that eqn(35) satisfies the necessary coincidence limits of a world function i.e. $\Omega_{\rm schw}]=[\nabla_{a}\Omega_{\rm schw}]=0\ ,\ [\nabla_{a}\nabla_{b}\Omega_{\rm schw}]=g_{ab}=\text{diag}(-1,(t/t_{0})^{2/3},(t/t_{0})^{2/3},(t/t_{0})^{-1/3})$ and gives a self consistent series expansion of $\Omega$ in Schwarzschild case. Furthermore, we calculated the van Vleck determinant for Kasner limit of Schwarzschild spacetime, which shows different scaling behaviour in singular and coincidence limits.
$\Delta_{\text{schw}}$ $\simeq$ $1+\mathcal{O}(\varepsilon^{4})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ :\ \textrm{Coincidence limit}$ (37)
$\Delta_{\text{schw}}$ $\simeq$ $\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}+\mathcal{O}(T^{1/3})\ \ \ \ :\ \ \textrm{Singularity limit}$ (38)
@@ III Geodesic and causal structure near singularity
The equi-geodesic surfaces are defined as surfaces of constant bi-scalar $\Omega(p,\mathscr{P})$ about a fiducial spacetime point. These surfaces constitute a fundamental geometric structure which has been extensively explored in Stargen and Kothawala (2015)Kothawala (2013). Let us now consider the global behaviour of these equi-geodesic surface $\Sigma(\kappa)$ defined by $\{p\ |\ \Omega(p,P)=-\kappa\ ,\ p\in\mathcal{M}\}$ in a FLRW spacetime with one of the points of world function $\Omega$ anchored at $P=(T,X,Y,Z)$. The equations for $\Sigma(\kappa)$ and lightcones are given by,
@@
We plot the set of equations (39) in the vicinity of the singularity $\mathscr{P}$ about different points $P$ translated in time to study the causal structure near the singularity. Since FLRW spacetime is homogeneous, we consider $X,Y,Z=0$ without loss of generality. The lightcones flatten out in the FLRW case (see fig 3) but still remain asymptotes to the time-like hypersurfaces. We find a contrasting behavior for Schwarzschild geometry, the right figure in 3 shows light cones drawn at different points $P$ along the radial infalling geodesic of a point particle in the vicinity of the Schwarzschild singularity. The light cones begin to shrink and converge as we approach the Schwarzschild singularity along the contracting dimension. The Schwarzschild metric in Kasner form is given by,
@@
$\tilde{K}=K-K_{0}\simeq\frac{\ddot{a}(t)}{a(t)}\varepsilon+\mathcal{O}(\varepsilon^{2})$ (46)
@@
where $a_{i}$ are defined as in eq.(32). The behaviour of lightcones in this case should be contrasted with the Kasner case for Schwarzschild singularity (see 4 and right half of Fig. 3) where lightcones close up along the contracting dimension (here $z$), reducing the region of spacetime that remains in causal contact for the timelike observer as singularity is approached.

### v2

@@ I Introduction
An important first step in understanding the role of quantum effects in regions of strong gravitational fields is to understand the behavior of geodesic flows - behavior of geodesics emanating from a single point - in vicinity of the singularity. This is so because many important features of classical and quantum dynamics depend on this behavior; see, for instance, Gutzwiller (1990) for an extensive textbook discussion on this point. While the structure of these flows is well understood in geodesically convex neighborhoods of regular spacetime points, not much is known about their behavior in regions containing curvature singularities. This is to be expected since covariant Taylor expansions that work well in the former case breaks down in the latter, and hence the analysis of geodesic flows directly in terms of curvature related quantities can not be done for generic singularities. We derive here, for the first time, explicit behavior of Synge’s world function $\Omega(x,y)$ and van Vleck determinant $\Delta(x,y)$ - two important bi-scalars characterizing geodesic flows - in the vicinity of FLRW and Kasner-like singularities. In the process, we improve upon an old result by Buchdahl [Buchdahl (1972)Buchdahl and Warner (1980) who calculated the series expansion of $\sqrt{|\Omega|}$” in FLRW and obtained a series whose terms can be shown to diverge as one approaches the singularity at $t\to 0$. This obviously un-physical result, as we shall see, is an artifact of expanding a wrong function (and manifests, in fact, even in the Minkowski limit!), and was presumably not noticed by Buchdahl. Here, we focus instead on $\Omega(x,y)$ itself, which we take to be the only meaningful quantity.
@@ I.1 Synge’s world function and van Vleck determinant
Synge’s world function and van Vleck determinant are two important bi-scalars that characterize measurements in curved spacetimes, and also appear ubiquitously in quantum field theory. Two-point (and higher) correlators of quantum fields in curved spacetime have a universal expansion in the world function near the coincidence limit. These expansions form the backbone of the point-splitting program for regularization in curved spacetime, and has been developed and employed extensively by de Witt, Cristensen, Parker and others in deriving covariant expansions of effective lagrangians in curved spacetimes, covariant formulation of path integral etc. The use of the world function as a fundamental object to study spacetime dates back to Synge, where it is defined operationally as half the square of the geodesic distance between the two spacetime points (denoted by $\Omega(x,x^{\prime})$). The spacetime metric $g_{ab}(x)$ can be constructed as the coincidence limit of $\nabla_{a}\nabla_{b}\Omega(x,y)$, i.e.
@@
Similarly, higher order coincidences gives us information about the spacetime curvature. Thus world function can be used to study tensor expansions on a given spacetime manifold in a covariant manner. The only hindrance to it’s use (apart from possible existence of multiple geodesics connecting two points) is the non-existence of any closed-form expression for geodesic interval as a function of spacetime coordinates even in highly symmetric spacetimes. In fact, to the best of our knowledge, closed form expressions are known only in maximally symmetric and plane-wave spacetimes Hari and Kothawala (2021),Harte and Drivas (2012). It would surprise the in-cognoscenti that the simplest and most widely encountered spacetimes such as Schwarzschild and FLRW have no closed form expression for $\Omega$. In this work, we derive a systematic representation for the world function in FLRW and Bianchi spacetimes, in a form that is applicable near singularities in these spacetimes.
@@
Another important bi-tensor of central importance that can be computed from the world function directly is the van Vleck determinant. It appears at various places, like fluctuation factor in semi-classical approximation to path integrals in curved spacetime, heat kernel expansions in curved spacetime, reconstruction of spacetime from two-point correlators, etc. Given a geodesic with endpoints $x$ and $z$, the van Vleck determinant is defined as,
@@
Intuitively, $\Delta(x,z)$ gives a quantitative measure of density of geodesics emanating from a given point $z$ and measured at points $x$ a fixed geodesic distance from $z$; this is depicted in Fig. 1.
@@ I.2 Summary of results
Since our analysis will be somewhat involved, we summarize our main results in Table 1, which gives the coincidence and singularity limits for the various biscalars in FLRW and Bianchi Type I spacetime. The reader is requested to refer to the relevant section(s) for definitions of various quantities in the table, which we have excluded here to avoid clutter. We will also present an infinite series representation for Synge’s world function for FLRW and Kasner (Schwarzschild) spacetime. The limits of $\Delta$ and $\Box\Delta^{1/2}$ and $K=\textrm{Tr}(K_{ab})$ are used to study the geodesic and causal structure near the curvature singularity. The structure of light cones is illustrated in Fig. 3 and Fig. 4 by plotting null geodesics along with the $\Omega=$ constant surfaces for different (fixed) points in the neighborhood of FLRW and Kasner (Schwarzschild) singularity.
@@
$\Delta(t,T)$ $\textrm{matter}:1+\frac{\varepsilon^{2}}{9T^{2}}+\cdots$ $\textrm{radiation}:1+\frac{\varepsilon^{2}}{8T^{2}}+\cdots$ $\textrm{matter}:-\frac{\ell^{2}t}{972\ T^{5/3}t_{0}^{4/3}}-\frac{\ell^{2}t}{243\ T^{4/3}t_{0}^{4/3}}$ $\ \ \ \ \ \ \ \ \ \ \ +\frac{t}{27T}\left(1-\frac{\ell^{2}(t/t_{0}^{4})^{1/3}}{4t}\right)$ $\textrm{radiation}:\frac{\ell^{2}t^{3/2}(5-2\log(t/T))}{4\ T^{5/2}t_{0}(\log(t/T))^{6}}$ $\ \ \ \ \ \ \ \ \ +\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}$ $1+O(\varepsilon^{4})$ $-\frac{5\ell_{\perp}^{2}(t/t_{0}^{4})^{1/3}}{486\ T}$ $-\frac{10\ell_{\perp}^{2}}{243\ T^{2/3}\ t_{0}^{4/3}}$ $+\frac{25t_{0}^{2/3}(z-Z)^{2}}{286\ t^{7/3}T^{1/3}\ t_{0}^{4/3}}$ $+\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}$ $-\frac{25\ell_{\perp}^{2}}{243\ (t\ t_{0}^{4})^{1/3}T^{1/3}}$
$\Box\Delta^{1/2}(t,T)$ $K(t,T)$ $\textrm{matter}:\frac{2}{9T^{2}}$ $\textrm{radiation}:\ \ 0$ $\frac{3}{\varepsilon}+\frac{\ddot{a}}{a}\varepsilon$ $\textrm{matter}:\frac{2267}{2187\sqrt{21}t^{2}}$ $\textrm{radiation}:\ \ \ 0$ $\frac{3q}{t}$ $\frac{3}{\varepsilon}-\frac{8\varepsilon^{3}}{1215T^{4}}$ $\frac{8}{3t}$
@@
The details of the results presented in Table 1 will be discussed in the rest of the article. However, from the results summarized in the table above, we can immediately compare our series representation for $\Omega(t,T)$ for FLRW spacetime, with that given by Buchdahl for $\sqrt{|\Omega|}$, in the limit of small $\varepsilon$, where $\varepsilon=t-T$,
$\Omega(t,T)\simeq-\frac{\varepsilon^{2}}{2}+\frac{\ell^{2}}{2}\Biggl(a(T)^{2}-a(T)\dot{a}(T)\ \varepsilon+\frac{1}{3}a(T)\dot{a}(T)\ \varepsilon^{2}\Biggl)+\frac{\ell^{4}}{24}\Biggl(a(T)^{2}\dot{a}(T)^{2}-a(T)\dot{a}(T)(\dot{a}(T)^{2}+a(T)\ddot{a}(T))\varepsilon\Biggl)$ (3)
On the other hand, the so called Hamilton/characteristic function $\sqrt{|\Omega|}$” defined by Buchdahl in [Buchdahl (1972), gives divergences in coincidence limit at $\mathcal{O}(\ell^{4})$:
$\sqrt{|\Omega|}\simeq-\frac{a(T)^{4}}{\varepsilon^{3}}+\textrm{less singular terms}$ (4)
Our series representation (3) is non-divergent both in the singularity limit $a(T)\to 0$ as well as in the coincidence limit $\varepsilon\to 0$, Thus $\Omega(t,T)$ serves as a more natural biscalar to study geometric properties of spacetime.
@@ II.1 FLRW spacetime
The Eq.(11) is a non-perturbative closed-form expression in terms of the geodesic length $\ell$ in geodesic distance on the three-space of constant curvature. The complication of further simplification arises because of the time-dependent scale factor $a(t)$, which is kept open for generality. As another self-consistency check, it can be verified that Eq.(11) satisfies the HJ equation up to order $\mathcal{O}(\ell^{4})$ with;
@@
$\Delta(t,T)=\frac{1}{a(t)^{3}a(T)^{3}F(t,T)^{3}}+\frac{20F(t,T)^{3}Q(t,T)-F^{(1,1)}(t,T)}{2a(t)^{3}a(T)^{3}F(t,T)^{5}}\ell^{2}+\mathcal{O}(\ell^{4})$ (13)
@@
where $I_{n}(a_{i})$ is defined by Eq.(33). We now consider the limiting behaviour of $\Delta(t,T)$ along a timelike geodesic in coincidence and singularity limit in matter and radiation dominated FLRW spacetime. The expansion in $\epsilon=t-T<<T$ (Coincidence limit) and (Singularity limit) $T<<t,T\to 0$ are obtained as follows,
@@
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $=$ $1+\frac{\epsilon^{2}}{9T^{2}}+\mathcal{O}(\epsilon^{3})$ (17)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $=$ $-\frac{\ell^{2}t}{972\ T^{5/3}t_{0}^{4/3}}-\frac{\ell^{2}t}{243\ T^{4/3}t_{0}^{4/3}}+\frac{t}{27T}\left(1-\frac{\ell^{2}(t/t_{0}^{4})^{1/3}}{4t}\right)+$... (18)
@@
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $=$ $1+\frac{\epsilon^{2}}{8T^{2}}+\mathcal{O}(\epsilon^{3})$ (19)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $=$ $\frac{\ell^{2}t^{3/2}(5-2\log(t/T))}{4\ T^{5/2}t_{0}(\log(t/T))^{6}}+\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}+$... (20)
@@ II.2 Bianchi Type I singularities
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are integration constants. The Eq.(27) can be inverted using Lagrange-Good formula Abdesselam (2003) for multivariables to give a series expansion in (xX),(yY),(zZ)(x-X),(y-Y),(z-Z), where
@@
where $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $a,b=1,2,3$. The $\mathcal{O}((x_{i}-X_{i})^{4})$ expansion for $\Omega$ using Eq.(29) gives,
@@
Where $\ell^{2}_{\perp}=(x-X)^{2}+(y-Y)^{2}$. It is easy to verify that Eq.(37) satisfies the necessary coincidence limits of a world function i.e. $\Omega_{\rm schw}]=[\nabla_{a}\Omega_{\rm schw}]=0\ ,\ [\nabla_{a}\nabla_{b}\Omega_{\rm schw}]=g_{ab}=\text{diag}(-1,(t/t_{0})^{2/3},(t/t_{0})^{2/3},(t/t_{0})^{-1/3})$ and gives a self consistent series expansion of $\Omega$ in Schwarzschild case. Furthermore, we calculated the van Vleck determinant for Kasner limit of Schwarzschild spacetime, which shows different scaling behaviour in singular and coincidence limits.
$\Delta_{\text{schw}}$ $=$ $1+\mathcal{O}(\varepsilon^{4})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ :\ \textrm{Coincidence limit}$ (39)
$\Delta_{\text{schw}}$ $=$ $-\frac{5\ell_{\perp}^{2}(t/t_{0}^{4})^{1/3}}{486\ T}-\frac{10\ell_{\perp}^{2}}{243\ T^{2/3}\ t_{0}^{4/3}}+\frac{25t_{0}^{2/3}(z-Z)^{2}}{286\ t^{7/3}T^{1/3}\ t_{0}^{4/3}}$ (41)
$\ \ -\frac{25\ell_{\perp}^{2}}{243\ (t\ t_{0}^{4})^{1/3}T^{1/3}}+\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}+...\ \ \ \ :\ \ \textrm{Singularity limit}$
@@ III Geodesic and causal structure near singularity
The equi-geodesic surfaces are defined as surfaces $\Omega(p,\mathscr{P})=$ const. about a spacetime point $\mathscr{P}$. These surfaces constitute a fundamental geometric structure which has been extensively explored in Stargen and Kothawala (2015)Kothawala (2013). We will now consider the behaviour of these equi-geodesic surface $\Sigma(\kappa)$ defined by $\{p\ |\ \Omega(p,P)=-\kappa\ ,\ p\in\mathcal{M}\}$ in a FLRW spacetime with one of the points of world function $\Omega$ anchored at $P=(T,X,Y,Z)$. The equations for $\Sigma(\kappa)$ and lightcones are given by,
@@
We plot the set of Eqs. (42) in the vicinity of the singularity $\mathscr{P}$ about different points $P$ translated in time to study the causal structure near the singularity. Since FLRW spacetime is homogeneous, we consider $X,Y,Z=0$ without loss of generality. The light cones flatten out in the FLRW case as seen from Fig. 3 (left). We find a contrasting behavior for Schwarzschild geometry, as is evident from Fig. 3 (right), which shows light cones drawn at different points $P$ along the radial infalling geodesic of a point particle in the vicinity of the Schwarzschild singularity. The light cones begin to shrink and converge as we approach the Schwarzschild singularity along the contracting dimension. The Schwarzschild metric in Kasner form is given by,
@@
$\tilde{K}=K-K_{0}=\frac{\ddot{a}(t)}{a(t)}\varepsilon+\mathcal{O}(\varepsilon^{2})$ (49)
@@
where $a_{i}$ are defined as in Eq.(34). The behaviour of lightcones in this case should be contrasted with the Kasner case for Schwarzschild singularity (see Fig. 4 and right half of Fig. 3) where lightcones close up along the contracting dimension (here $z$), reducing the region of spacetime that remains in causal contact for the timelike observer as singularity is approached.